I drag a mass m=21 kg in a straight line, along a horizontal surface, a distance D=23 m . I drag it at constant speed v=0.90 m.s−1 in a straight line using a horizontal force. The coefficients of friction are \mu_{s} = 1.2μ
s

=1.2 and \mu_{k} = 1.1μ
k

=1.1. How much work do I do? W =W= _____ J
1
Expert's answer
2020-09-10T11:59:09-0400

The speed is said to be constant, so the acceleration is equal to zero. Therefore, the sum of the horizontal force and the friction force is zero. Let us calculate the friction force. It is equal to "\mu_k N = \mu_kmg = 1.1\cdot 21\cdot 9.8 \approx 226.4\, \mathrm{N}."

We should take into account the coefficient of kinetic friction, because the mass is said to move with constant velocity, we don't consider the beginning of the motion.

The modulus of work of the friction force is equal to the work of the horizontal force and can be obtained as "\mu_kmgD = 226.4\cdot 23 \approx 5207\,\mathrm{J}."